Please help!
My house is 10 miles from the airport. A plane took off from the airport and when it was directly over my house it had traveled 12 miles. How many miles high is the airplane when it is directly over my house?

Found 2 solutions by mananth, scott8148:
Answer by mananth(16946)

the ground distance to airport = 10 miles.
when plane is above your house the plane has flown 12 miles
.
let x be the height of plane above your house.
house, the plane and the airport form a right triangle.
ground distance and height are the legs.
12^=10^2+x^2
x^2=12^2-10^2
x^2=144-100
x^2=44
x= 6.6 miles above your house
...
m.ananth@hotmail.ca

Answer by scott8148(6628)

the plane's path is the hypotenuse of a right triangle

the distance to the airport is one leg , and the height of the plane is the other

by Pythagoras ___ 10^2 + h^2 = 12^2